# Eliminating Biased Length Reliance of Direct Preference Optimization via Down-Sampled KL Divergence

### v2

@@ Abstract
Junru Lu1∗, Jiazheng Li2thanks: Equal Contributions by Junru Lu and Jiazheng Li, Siyu An3, Meng Zhao3, Yulan He2, Di Yin3, Xing Sun3 1University of Warwick, 2King’s College London, 3Tecent YouTu Lab 1junru.lu@warwick.ac.uk, 2{jiazheng.li, yulan.he}@kcl.ac.uk 3{siyuan, alexmzhao, endymecyyin, winfredsun}@tencent.com
@@ 3.3 Debiased KL Divergence
where $\bm{T}_{m}$ equals to the minimum token length of $(\bm{T}_{w},\bm{T}_{l})$, and $\bm{y}^{t}$ is down-sampled from all tokens $\{\bm{y}^{T}\}$ uniformly. Eq. 10 is consistent with the corresponding reward term in Figure 1, as shown in the middle. In addition, we provide ablation studies of sampling randomness in Appendix G.
@@ 5 Experimental Results
In this section, following above designs, we first report the group experiments of length reliance (§ 5.1), then present comparison studies against strong baselines (§ 5.2). We discuss quantitative results in the main body. We leave more ablation studies and case analysis in Appendix G,  H, and  J.
Figure 3: Trends of DPO’s biased implicit reward (Eq. 7), given different sets. Debiased rewards: App. F.
@@
Table 1: Performance of models in Figure 3.
5.1 Group study of length reliance
The “-all” set refers to training with original UltraFeedback, which mix “-long” and “-short” data. The “-long” subset provides overestimated rewards and therefore causes performance degradation. However, since statistically, the chosen response does being longer than the rejected response Park et al. (2024), the training trend of the “-long” subset is similar to the “-all” full set. On the contrary, the “-short” subset completely erases the distinctive feature of length, hoping that the model will perform comparative learning based entirely on content quality. However, the biased DPO completely underestimate the reward, thus causing collapses. Appendix F shows our debaised rewards.
@@ 5.2.2 Study on HH-RLHF & TL;DR
For study on HH-RLHF and TL;DR, we report the win rate against the SFT basis, and the average generated token length of all methods in Table 3.
@@
SamPO achieves the best performance on TL;DR. In terms of TL;DR, SamPO and Iterative SamPO show the highest win rates, with 65.71% and 73.58% respectively. DPO and Length-normed DPO also perform well, achieving win rates of 60.98% and 58.13% respectively. Iterative DPO reaches the best while using longer answers than Iterative SamPO. In contrast, SimPO has the lowest win rate at 33.33%, indicating that it is less effective on the TL;DR dataset.
Over-simplification by SimPO. In fact, on HH-RLHF, we notice many of the outputs from SimPO are overly simplified, often omitting necessary content and resulting in only 14.77 length of tokens on average. For example, a preferred response from HH-RLHF is “I’ll give you the links.”, whereas the SimPO response is simply “Sure!”. This suggests that while concise, the responses lack the necessary informativeness. In this scenario, we can see GPT-4 prefers over-simplified responses, which probably due to the binary setup of preference choice. Similarly, on TL;DR, SimPO produces the shortest responses (averagely 31.90 tokens). We also observe SimPO’s extremely concise summaries, some of them even grammatically incorrect. For example, SimPO outputs a summary without subject : “online flirt turns into legit relationship. Great chemistry. Age gap and distance issues. Need advice before final meetup before long trip abroad.”.
@@ 6 Conclusion
While our proposed method, SamPO, has shown promising results in mitigating verbosity and improving performance, there are several limitations and areas for future work:
  • Scalability. Although we test SamPO on three different LLMs, following the common practice with most DPO-related works. We agree that further experiments are needed to confirm its scalability and generalization across a broader range of much larger models (e.g., Qwen-110B Bai et al. (2023)).
  • Computational Overhead. The down-sampling approach introduces additional computational steps during training. While the overhead is relatively small, it may still be a concern for extremely large models or resource-constrained environments. Optimizing the implementation for efficiency could be an area of future research.
  • Human Evaluation. Our evaluations of open-ended benchmarks primarily rely on GPT-4 judgments. While it provides valuable insights, GPT-4 may contain potential bias and not fully capture the nuances of human preferences. Conducting human evaluations would offer a more accurate assessment of the quality and alignment of the generated responses.
@@ Acknowledgment
We thank Shiyue Xu for pointing out the error in Equation 5 in the previous draft.
@@ A.2 Gradients of Token-level DPO reward
$=\nabla_{\theta}\log\prod_{t=1}^{T_{w}}\pi(y_{w,t}|y_{w,<t},x)-\nabla_{\theta}\log\prod_{t=1}^{T_{l}}\pi(y_{l,t}|y_{l,<t},x)$
@@
$\mathcal{M}$ $=\nabla_{\theta}\log\prod_{t=1}^{T_{m}}\pi(y_{w,t}|y_{w,<t},x)-\nabla_{\theta}\log\prod_{t=1}^{T_{m}}\pi(y_{l,t}|y_{l,<t},x)$
@@ Appendix B Evaluation Details
For all conditional benchmarks, we use a stable and popular evaluation framework “lm-evaluation-harness” Gao et al. (2023b) [^3] Official tool page of lm-eval: https://github.com/EleutherAI/lm-evaluation-harness. As for open-ended benchmarks, we report specific evaluation templates for AlpacaEval2, HH-RLHF and TL;DR in Appendix K. Particularly, we use the official tool to evaluate AlpacaEval2444https://github.com/tatsu-lab/alpaca_eval/. The version of GPT-4 evaluator is all set as: gpt-4-turbo.
@@
Appendix C HyperParameters and Training Cost
Appendix D Preliminary Study of DPO & Variants
Appendix E Random K vs. Top K
@@ Appendix D Preliminary Study of DPO & Variants
GSM8K IFEval PiQA MMLU TruthfulQA Avg.
long 41.24 37.89 81.28 55.86 38.68 50.99
short 34.50 6.00 77.09 54.87 30.48 40.59
all 42.61 43.76 81.77 55.85 35.86 51.97
long* 42.61 38.01 81.18 55.86 36.11 50.75
short* 41.70 33.93 81.18 55.5 36.35 49.73
all* 42.68 44.12 81.28 55.8 40.15 52.81
Appendix F Are rewards debiased?
We add SamPO-based group experiments to Figure 3, resulting in the new Figure 6. All the models fine-tuned in this figure is based on the same Tulu2-13B-SFT model. It can be seen that DPO leads to biased rewards, especially the collapse one of “-short” subset, while SamPO debaises former rewards. We can observe debiased positive rewards on the “-short” set. And the debaised rewards of “-all” set grow to a high peak at 300 steps.
Such debiased rewards result in significant U-turn reversal and further improvements. As reported in Table 7, SamPO manages to eliminate collapse on the “-short” set, where we record a normal average benchmark score similar to the “-long” set, improving the DPO’s result by 9.2%. Thanks to the regularization of those “short” data, the “-all” set that mix both “long” and “short” data, achieves the best performance up to 52.81 on average.
@@ Appendix F Are rewards debiased?
Appendix G Influence of Different Random Seed
Appendix H Influence of Different Beta in Eq. 1
@@ Appendix G Influence of Different Random Seed
MRC Logical Reasoning RolePlay Avg.
SFT Base 81.25 69.52 59.12 69.96
w/ DPO 85.33 73.25 57.41 72.00
w/ SamPO 87.50 83.57 63.61 78.23
@@
Appendix I Human Evaluation of SamPO
In addition to the aforementioned automated evaluation, we further conduct large-scale human evaluation to study the effectiveness of the SamPO algorithm when applied to super large LLM (e.g, over 50B). We use an LLM fine-tuned based on Qwen1.5-72B Bai et al. (2023) as a starting point and fine-tune it for 1 epoch using the proposed SamPO method. The training data is a general preference dataset of around 480,000 samples.
We report the results of human evaluation in Table 8, covering three most popular scenarios: general Machine Reading Comprehension (MRC), logical reasoning (e.g., math or logic questions), and open domain dialogues in role-play settings. We have hired a 30-person annotation team, each of whom has at least a bachelor’s degree or above. And each test scenario contains 500 to 1,000 carefully crafted challenging instances, which then cross-labeled by multiple professional annotators.
Our scoring criteria are relatively simple, distinguishing only between incorrect and acceptable responses, rather than employing a more fine-grained scoring system (such as a 1 to 10 point scale). We observe that SamPO significantly outperforms both the SFT Base and DPO methods on all three tasks.
Appendix J Case Analysis
In this section, we introduce exemplars to analyze the qualitative outputs from different approaches. We analyze two cases for each test set: AlpacaEval2 (§ J.1), HH-RLHF (§ J.2), and TL;DR (§ J.3), in each section respectively.
J.1 Case analysis on AlpacaEval2
@@ Appendix I Human Evaluation of SamPO
J.2 Case analysis on HH-RLHF
@@
Appendix K GPT-4 Judged Winner Templates

### v3

@@ Abstract
Junru Lu1∗,3, Jiazheng Li2thanks: Equal Contribution., Siyu An3, Meng Zhao3, Yulan He1,2,4, Di Yin3, Xing Sun3 1University of Warwick  2King’s College London 3Tencent YouTu Lab  4The Alan Turing Institute junru.lu@warwick.ac.uk, {jiazheng.li, yulan.he}@kcl.ac.uk {siyuan, alexmzhao, endymecyyin, winfredsun}@tencent.com
@@ 3.3 Debiased KL Divergence
where $\bm{T}_{m}$ equals to the minimum token length of $(\bm{T}_{w},\bm{T}_{l})$, and $\bm{y}^{t}$ is down-sampled from all tokens $\{\bm{y}^{T}\}$ uniformly. Eq. 10 is consistent with the corresponding reward term in Figure 1, as shown in the middle. In addition, we provide ablation studies of sampling randomness in Appendix E.
@@ 5 Experimental Results
In this section, following above designs, we first report the group experiments of length reliance (§ 5.1), then present comparison studies against strong baselines (§ 5.2). We discuss quantitative results in the main body. We leave more ablation studies and case analysis in Appendix E,  F, and  H.
@@
long* 42.61 38.01 81.18 55.86 36.11 50.75
short* 41.70 33.93 81.18 55.5 36.35 49.73
all* 42.68 44.12 81.28 55.8 40.15 52.81
@@
5.1 Group study of length reliance
The “-all” set refers to training with original UltraFeedback, which mix “-long” and “-short” data. The “-long” subset provides overestimated rewards and therefore causes performance degradation. However, since statistically, the chosen response does being longer than the rejected response Park et al. (2024), the training trend of the “-long” subset is similar to the “-all” full set. On the contrary, the “-short” subset completely erases the distinctive feature of length, hoping that the model will perform comparative learning based on content quality. However, the biased DPO completely underestimate the reward, thus causing collapses.
Yet, our SamPO presents debaised rewards. We can observe debiased positive rewards on the “-short” set. And the debaised rewards of “-all” set grow to a high peak at 300 steps. Such debiased rewards result in significant U-turn reversal and further improvements. As shown in Table 1, SamPO manages to eliminate collapse on the “-short” set, where we record a normal average benchmark score similar to the “-long” set, improving the score by 9.2%. Thanks to the regularization of those “short” data, the “-all” set that mix both “long” and “short” data, achieves the best score up to 52.81 on average.
@@ 5.2.2 Study on HH-RLHF & TL;DR
As for HH-RLHF and TL;DR, we utilize Pythia-2.8B for all experiments. Given that Pythia has not been specifically trained for instructional tasks, we initiate our process with one epoch of SFT on the chosen response, following DPO’s setup. Subsequently, we conduct preference optimization using SamPO alongside various baseline methods. Following previous literature Rafailov et al. (2023); Park et al. (2024), GPT-4 served as the proxy for human preference. We report the win rate against the SFT basis, and the average generated token length of all methods in Table 3.
@@
SamPO achieves the best performance on TL;DR. In terms of TL;DR, SamPO and Iterative SamPO show the highest win rates, with 65.71% and 73.58% respectively, significantly outperforming all other methods. DPO and Length-normed DPO also perform well, achieving win rates of 60.98% and 58.13% respectively. Iterative DPO reaches the best while using longer answers than Iterative SamPO. In contrast, SimPO has the lowest win rate at 33.33%, indicating that it is less effective on the TL;DR dataset.
Over-simplification by SimPO. In fact, on HH-RLHF, we notice many of the outputs from SimPO are overly simplified, often omitting necessary content and resulting in only 14.77 length of tokens on average. For example, a preferred response from HH-RLHF is “I’ll give you the links.”, whereas the SimPO response is simply “Sure!”. This suggests that while concise, the responses lack the necessary informativeness. In this scenario, we can see GPT-4 prefers over-simplified responses, which probably due to the binary setup of preference choice. Similarly, on TL;DR, SimPO produces the shortest responses (averagely 31.90 tokens). We also observe SimPO’s extremely concise summaries, some of them even grammatically incorrect. For example, a preferred summary from the TL;DR is “I [20M] met a great girl [16F] online who lives in the same city. Problems are: she’s moving away, I want to meet her, and the obvious age gap.”, while SimPO outputs a shorter summary without subject and capitalizes fist letter: “online flirt turns into legit relationship. Great chemistry. Age gap and distance issues. Need advice before final meetup before long trip abroad.”.
5.2.3 Human Evaluation of SamPO
In addition to the aforementioned automated evaluation, we further conduct large-scale human evaluation to study the effectiveness of the SamPO algorithm when applied to super large LLM (e.g, over 50B). We use an LLM fine-tuned based on Qwen1.5-72B Bai et al. (2023) as a starting point and fine-tune it for 1 epoch using the proposed SamPO method. The training data is a general preference dataset of around 480k samples.
We report the results of human evaluation in Table 4, covering three most popular scenarios: general Machine Reading Comprehension (MRC), logical reasoning (e.g., math or logic questions), and open domain dialogues in role-play settings. We have hired a 30-person annotation team, each of whom has at least a bachelor’s degree or above. And each test scenario contains 500 to 1k carefully crafted challenging instances, which then cross-labeled by multiple professional annotators. Our scoring criteria are relatively simple, distinguishing only between incorrect and acceptable responses. We observe that SamPO significantly outperforms both the SFT Base and DPO method on all tasks.
@@ 6 Conclusion
While our proposed method, SamPO, has shown promising results in mitigating verbosity and improving performance, several limitations remain:
  • Scalability. Although we test SamPO on different LLMs, including one super large LLM (Qwen1.5-72B-Instruct). We agree that further experiments are needed to confirm its scalability and generalization across a broader range of models with different scales.
  • Computational Overhead. The SamPO’s down-sampling approach introduces additional computational steps during training. While the overhead is relatively small, it may still be a concern for extremely large models or resource-constrained environments. Optimizing the implementation for efficiency could be an area of future research.
  • Human Evaluation. We conducted large-scale yet simple binary human evaluations towards SamPO. Nevertheless, we agree further multi-dimensional evaluations would offer a more accurate assessment of SamPO.
@@ Acknowledgment
We thank Shiyue Xu for correcting the error in Equation 5 in the previous draft [^3] https://github.com/LuJunru/SamPO/issues/1. This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2) and Innovate UK through its Accelerating Trustworthy AI Collaborative R&D funding (grant no. 10093055).
@@ A.2 Gradients of Token-level DPO reward
$=\nabla_{\theta}\log\prod_{t=1}^{T_{w}}\pi(y_{w,t}|y_{w,<t},x)$
$\quad-\nabla_{\theta}\log\prod_{t=1}^{T_{l}}\pi(y_{l,t}|y_{l,<t},x)$
@@
$\mathcal{M}$ $=\nabla_{\theta}\log\prod_{t=1}^{T_{m}}\pi(y_{w,t}|y_{w,<t},x)$
$-\nabla_{\theta}\log\prod_{t=1}^{T_{m}}\pi(y_{l,t}|y_{l,<t},x)$
@@ Appendix B Evaluation Details
For all conditional benchmarks, we use a stable and popular evaluation framework “lm-evaluation-harness” Gao et al. (2023b) [^4] Official tool page of lm-eval: https://github.com/EleutherAI/lm-evaluation-harness. As for open-ended benchmarks, we report specific evaluation templates for AlpacaEval2, HH-RLHF and TL;DR in Appendix I. Particularly, we use the official tool to evaluate AlpacaEval2555https://github.com/tatsu-lab/alpaca_eval/. The version of GPT-4 evaluator is all set as: gpt-4-turbo.
@@
Appendix C HyperParameters and Training Cost
Appendix D Preliminary Study of DPO & Variants
Appendix E Influence of Different Random Seed
@@
Appendix F Influence of Different Beta in Eq. 1
@@
Appendix G Random K vs. Top K
@@
Appendix H Case Analysis
In this section, we introduce exemplars to analyze the qualitative outputs from different approaches. We analyze two cases for each test set: AlpacaEval2 (§ H.1), HH-RLHF (§ H.2), and TL;DR (§ H.3), in each section respectively.
H.1 Case analysis on AlpacaEval2
@@
H.2 Case analysis on HH-RLHF
@@ H.3 Case analysis on TL;DR
Appendix I GPT-4 Judged Winner Templates